OpenStudy (anonymous):

Find the additive and multipliactive inverse for each number: 3A. 1.25

OpenStudy (shane_b):

Do you know what additive and multiplicative inverse means? We can start there...

OpenStudy (anonymous):

I know what it means a little bit but not much. We just started this in our Algerba II

OpenStudy (shane_b):

Additive inverse is simply the number that when added to the other number equals 0. So \[1.25+? = 0\]

OpenStudy (anonymous):

25 + 0 = 0? Is thar right?

OpenStudy (anonymous):

That**

OpenStudy (shane_b):

First, let me clear something up: "3A. 1.25" So the problem is 3A and the value given is 1.25 right?

OpenStudy (anonymous):

Yepp

OpenStudy (shane_b):

Ok...so what could you add to 1.25 which would result in a value of 0?

OpenStudy (anonymous):

0?

OpenStudy (shane_b):

1.25 + 0 = 1.25 That's not it :)

OpenStudy (shane_b):

Since 1.25 is positive...you should see that you must add a negative number to make it 0. Therefore, the answer would be -1.25:\[1.25+(-1.25)=0\]That's a true equation.

OpenStudy (anonymous):

Right so it would be a -1.25 as a answer?

OpenStudy (shane_b):

Yes

OpenStudy (anonymous):

Alright so basically just switching a positve to a negative answer.

OpenStudy (shane_b):

Maybe... If you're original number was -1.25...the additive inverse would actually be +1.25

OpenStudy (shane_b):

If you're ok with everything above...we can discuss multiplicative inverses

OpenStudy (anonymous):

Yeah thats alright so for 3A it would be -1.25?

OpenStudy (shane_b):

Yes, that's additive inverse answer.

OpenStudy (shane_b):

You still need to do the multiplicative inverse also. Multiplicative inverse is simply defined as the number that when multiplied times the original number result in an answer of 1. In essence, it's going to be the reciprocal of the original value. \[(1.25)(\frac{1}{1.25})=1\]So the multiplicative inverse of 1.25 is 1/1.25. You could also express this (as well as the additive inverse as a fraction which may be what your teacher wants:\[1.25=\frac{5}{4}\]So the additive inverse would be :\[-\frac{5}{4}\] and the multiplicative inverse would be \[\frac{4}{5}\]